# USED: Universal Speaker Extraction and Diarization

### v2

@@ V-E Ablation Study of USED Models Trained on Single Task
V-F Investigation of the Relationship between Mult-Task Interaction Module and SAD Loss
@@
In this section, we analyze the performance of the USED-F model using different weights for each scenario of SAD loss with or without the multi-task interaction module as illustrated in Table IV. For the experiments without multi-task interaction module, results demonstrate that when increasing the weights $\gamma$ and $\delta$ for QS and QQ scenarios, we can get a lower power value, but the SI-SDRi are reduced from 12.02 dB to 10.48 dB for SS scenario and from 9.58 dB to 7.42 dB for SQ scenario. Regarding the performance of speaker diarization, DER results increase when larger values for the weights $\gamma$ and $\delta$ of QS and QQ scenarios are used. These observations indicate that the model performance under different scenarios is sensitive to the loss weights under different scenarios. With the help of the multi-task interaction module, we can keep the performance for speaker extraction under SS and SQ scenarios and speaker diarization but achieve a much lower power, which decreases from -1.93 dB/s to -21.04 dB/s.
@@ VI Conclusion
We propose USED, a unified network for universal speaker extraction and diarization, which integrates speaker extraction and diarization for managing speech mixture with varying overlap ratios. We design an embedding assignment module to support producing results for a variable number of speakers based on speech references and enhance the model’s robustness. Additionally, a multi-task interaction module is designed to leverage information from both speaker extraction and diarization tasks. Experimental results demonstrate significant improvements in both highly and sparsely overlapped speech scenarios for the speaker extraction and speaker diarization tasks.

### v3

@@ V-D Investigation of USED Models for Generating Results of Different Numbers of Speakers During Inference
V-E Ablation Study: Assessing Single-Task Training for the USED Model
@@
V-F Investigation of the Relationship between the Multi-Task Interaction Module and the SAD Loss
TABLE VII: Model Comparison in Terms of Parameters and MACs for the TS-VAD model, the SpEx+ model and the USED model. “Params” refers to the number of model parameters, which is counted in millions (M). †The MACs of SpEx+ is only for extracting one speaker’s result.
Model Params MACs Total MACs
TS-VAD [14] 39.50 27.21 158.01
SpEx+ †[9] 16.35 43.60
USED-F 23.12 96.91 96.91
USED 23.65 119.54 119.54
@@ V-F Investigation of the Relationship between the Multi-Task Interaction Module and the SAD Loss
In this section, we analyze the performance of the USED-F model using different weights for each scenario of SAD loss with or without the multi-task interaction module as illustrated in Table V. For the experiments without the multi-task interaction module, results demonstrate that when increasing the weights $\alpha$ and $\beta$ for QQ and QS scenarios, we can get a lower power value, but the SI-SDRi are reduced from 12.02 dB to 10.48 dB for the SS scenario and from 9.58 dB to 7.42 dB for the SQ scenario. Regarding the performance of speaker diarization, DER results increase when larger values for the weights $\alpha$ and $\beta$ of QQ and QS scenarios are used. These observations indicate that the model performance is sensitive to the loss weights under different scenarios. With the help of the multi-task interaction module, we can keep the performance for speaker extraction under SS and SQ scenarios and speaker diarization and simultaneously achieve a much lower power under the QQ scenario, which decreases from -1.93 dB/s to -24.00 dB/s.
V-G Complexity Analysis
In this section, we conduct a complexity analysis of the USED model in comparison to the baseline models, TS-VAD and SpEx+, as shown in Table VII. To ensure a fair comparison, the duration of the speech input for all models is kept consistent, with both the speech mixture and the speech references set to 4 seconds. The experimental results indicate that, compared to individual baseline models, the USED model has more parameters than the SpEx+ model but fewer than the TS-VAD model. However, the computational complexity of the USED model is higher than both SpEx+ and TS-VAD.
TABLE VIII: Impact of hyperparameter choices for embedding assignment module on model performance.
$m$ Metrics $p_{sel}$ $p_{em}$ $p_{res}$
0.1 0.3 0.5 0.7 0.9 0.1 0.3 0.5 0.7 0.9 0.1 0.3 0.5 0.7 0.9
1 SI-SDRi $\uparrow$ 11.81 10.36 12.12 12.81 12.60 10.77 12.12 12.93 13.09 13.10 13.12 12.99 12.72 13.01 12.12
DER $\downarrow$ 6.80 8.79 6.47 5.83 6.19 6.67 6.47 5.87 5.63 5.62 5.75 5.83 6.21 5.76 6.47
2 SI-SDRi $\uparrow$ 13.28 13.30 13.58 13.62 12.81 13.68 13.58 13.83 13.56 13.42 13.77 13.76 13.83 13.72 13.58
DER $\downarrow$ 5.17 5.11 4.98 5.28 9.41 5.01 4.98 5.32 4.97 5.05 5.67 5.15 5.24 4.93 4.98
all SI-SDRi $\uparrow$ 14.71 14.45 14.29 14.07 13.10 14.25 14.29 14.31 14.26 14.27 14.31 14.28 14.35 14.25 14.29
DER $\downarrow$ 3.87 4.00 4.22 4.41 6.27 4.30 4.22 4.16 4.13 4.18 4.23 4.22 4.16 4.19 4.22
TABLE IX: Model performance regarding different loss weight values on the validation set.
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}=0.1$ $\lambda_{3}=0.5$ $\lambda_{3}=1.0$
SI-SDRi $\uparrow$ DER $\downarrow$ SI-SDRi $\uparrow$ DER $\downarrow$ SI-SDRi $\uparrow$ DER $\downarrow$
0.1 0.1 14.31 4.15 14.27 4.26 14.27 4.15
0.1 0.5 14.01 3.85 13.99 3.76 14.06 3.74
0.1 1.0 13.54 3.85 13.46 4.00 13.32 4.19
0.5 0.1 14.22 5.01 14.13 5.14 14.19 5.01
0.5 0.5 14.16 4.23 14.27 4.26 14.32 4.14
0.5 1.0 14.19 3.97 14.32 3.92 14.28 3.84
1.0 0.1 14.11 5.49 14.15 5.37 14.17 5.47
1.0 0.5 14.15 4.57 14.18 4.55 14.32 4.45
1.0 1.0 14.22 4.21 14.32 4.14 14.29 4.22
It is important to note that the USED model simultaneously addresses two tasks: speaker extraction and diarization, whereas TS-VAD and SpEx+ are designed for a single task, with SpEx+ specifically extracting the speech for only one speaker. Therefore, we calculate the ‘Total MACs’. ‘Total MACs’ refers to the total number of MACs required for the models to process a speech mixture containing three speakers and obtain the corresponding speaker extraction and diarization results for all three speakers. When comparing the ‘Total MACs’, the ‘Total MACs’ of the USED and USED-F models are significantly lower than that of the baseline models combined. Specifically, the USED-F model reduces the total MACs by approximately 39%, while the USED model achieves a reduction of around 24%.
V-H Loss Weight Selection
To investigate the impact of different loss weights on the model’s performance and explore how to select appropriate values for these weights, we conduct a series of experiments with three loss weights ($\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$), as illustrated in Table IX. We evaluate the model’s performance on the validation set of the LibriMix dataset with max mode by testing each loss weight with values of 0.1, 0.5, and 1.0. The experimental results indicate that there may be room for improvement in the model’s performance regarding the selection of loss weights, although we set the values of all three weights to 1.0 for other experiments. For example, when $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are 0.5, 1.0 and 0.5, respectively, the USED model performs slightly better on both the speaker extraction and diarization tasks.
V-I Impact on Model Performance of Hyperparameter Tuning for Embedding Assignment Module
In this section, we examine the influence of hyperparameter selection within the embedding assignment module on the overall performance of the model, as shown in Table VIII. Specifically, we focus on three key hyperparameters associated with the embedding assignment module, conducting our analysis using the validation set from the LibriMix dataset in max mode. We assess the impact of varying the hyperparameter values $p_{sel}$, $p_{em}$, and $p_{res}$, which are assigned values of 0.1, 0.3, 0.5, 0.7, and 0.9, while keeping the two other hyperparameters fixed to their default value, across multiple configurations where $m$ is set to 1, 2, and all.
The experimental results demonstrate that the influence of the three hyperparameters on model performance varies depending on the value of $m$. For instance, for hyperparameter $p_{sel}$, larger values lead to better results when $m$ equals 1, whereas smaller values yield better performance when $m$ equals all. After carefully evaluating the overall performance across different values of $m$, we selected a set of parameters with $p_{sel}=0.5$, $p_{em}=0.3$, and $p_{res}=0.9$ for relatively good performance.
@@ VI Conclusion
We propose USED, a unified network for universal speaker extraction and diarization, which integrates speaker extraction and diarization for managing speech mixtures with varying overlap ratios and variable number of speakers. We design an embedding assignment module to support producing results for a variable number of speakers based on speech references and enhance the model’s robustness. Additionally, a multi-task interaction module is designed to leverage information from both speaker extraction and diarization tasks. Experimental results demonstrate significant improvements in both highly and sparsely overlapped speech scenarios for the speaker extraction and speaker diarization tasks.